# EchoSight: Advancing Visual-Language Models with Wiki Knowledge

### v1

@@ Abstract
Knowledge-based Visual Question Answering (KVQA) tasks require answering questions about images using extensive background knowledge. Despite significant advancements, generative models often struggle with these tasks due to the limited integration of external knowledge. In this paper, we introduce EchoSight, a novel multimodal Retrieval-Augmented Generation (RAG) framework that enables large language models (LLMs) to answer visual questions requiring fine-grained encyclopedic knowledge. To strive for high-performing retrieval, EchoSight first searches wiki articles by using visual-only information, subsequently, these candidate articles are further reranked according to their relevance to the combined text-image query. This approach significantly improves the integration of multimodal knowledge, leading to enhanced retrieval outcomes and more accurate VQA responses. Our experimental results on the Encyclopedic VQA and InfoSeek datasets demonstrate that EchoSight establishes new state-of-the-art results in knowledge-based VQA, achieving an accuracy of 41.8% on Encyclopedic VQA and 31.3% on InfoSeek.
@@
Yibin Yan1,2           Weidi Xie1,2 1Shanghai Jiao Tong University        2Shanghai AI Lab https://go2heart.github.io/echosight
@@ 1 Introduction
Despite the recent accomplishments in developing Visual-language Models (VLMs) Achiam et al. (2023); Gemini Team et al. (2023); Abdin et al. (2024); Liu et al. (2024), knowledge-based VQA remains challenging. This complexity primarily stems from two aspects. (i) Existing VLMs struggle to adequately encode all essential knowledge, due to its limited model capacity, and infrequent inclusion of encyclopedic, long-tail information in their training data Mensink et al. (2023). (ii) The visual component of the questions often provides limited help in addressing the queries, as establishing a meaningful connection between entity knowledge and visual attributes can be difficult. For example, an image of a church alone does not reveal information about its construction date.
@@
Overall, we present three contributions: First, we propose a multimodal retrieval-augmented generation framework, termed as EchoSight, that enables LLMs to answer visual questions that require fine-grained encyclopedic knowledge; Second, we adopt a retrieval-and-reranking scheme to improve retrieval performance, specifically, it first searches images with visual-only information, and then conduct a fine-grained multimodal reranking on the candidates; Third, we conduct thorough experiments on both Encyclopedic VQA Mensink et al. (2023) and InfoSeek Chen et al. (2023) benchmarks, EchoSight demonstrates state-of-the-art performance on both benchmarks, significantly outperforming existing VLMs or other retrieval-augmented architectures.
@@ 2.2 Retrieval and Reranking
On the candidates side, we break each of the wiki articles into sections, with each section prefixed by the article’s title, for example, $a_{i}=\{\text{sec}_{1}^{i},\text{sec}_{2}^{i},\dots,\text{sec}_{p}^{i}\}$, and further encode them with Q-Former’s text encoder. We initialize the Q-Former with BLIP-2’s weights and fine-tunes with all parameters open except the visual encoder.
@@
$\text{sec}_{vl}=\operatorname*{arg\,max}_{\text{sec}\in\mathcal{E}_{v}}\left(\alpha\cdot S_{v}^{\text{sec}}+(1-\alpha)\cdot S_{r}^{\text{sec}}\right),$
where $\text{sec}_{vl}$ refers to the highest-ranked entry section produced by the reranker, $\alpha$ is a weight parameter that balances the visual similarity score $S_{v}^{\text{sec}}$ and the reranking score $S_{r}^{\text{sec}}$. Note that $S_{v}^{sec}$ is calculated in the visual-only search stage using the best-matched image from the wiki entry to which $sec$ belongs.
Reranker Training. Here, we implement hard negative sampling within a contrastive learning framework. Specifically, hard negative samples are specifically selected from examples that are visually similar yet contextually distinct, i.e., the initial visual-only retrieval efforts were unsuccessful. With such training, the reranker is thus forced to select the most relevant articles for the multimodal queries, enhancing the overall accuracy and effectiveness of the system Robinson et al. (2021).
@@
$\mathcal{L}=-\log\frac{\exp(\max_{1\leq i\leq N_{q}}\text{sim}(z_{m}^{i},z_{s}))}{\sum_{j=1}^{N}\exp(\max_{1\leq i\leq N_{q}}\text{sim}(z_{m}^{i},z_{s}^{j}))},$
where $z_{s}$ denotes the positive section embedding, and $N$ is the total number of samples including both positive and negatives.
@@ 3.3 Implementation Details
We adopt OneCycleLR Smith and Topin (2019) scheduler, with AdamW Loshchilov and Hutter (2018) optimizer. We use learning rate $10^{-4}$, batch size $6$, and the negative samples per example being $24$. For training the reranker module with 900K examples, 150K steps require 40 hours on 1 Nvidia A100 (80G).
@@ 3.4 Results
On Retrieval. The experiment results for the retrieval tasks across different configurations are detailed in Table 1 and Table 2. The CLIP I-T setting involves using CLIP for cross-modal similarity search, from the reference image to the Wikipedia article. The articles are represented as CLIP embedding of their title and descriptions. The ‘Google Lens’ refers to the approach used in Encyclopedic VQA Mensink et al. (2023), where Google Lens indexes billions of images from the Internet, not limited to Wikipedia, to find and return the most closely matching images along with an entity prediction. The best corresponding knowledge base entry identified by Google Lens is considered the result of its retrieval effort. Given its vast image index and capability to associate images with relevant entities, Google’s retrieval can be viewed as a upperbound in E-VQA retrieval.
From both tables, we can draw the observation that, our proposed reranking module has shown to significantly improve the retrieval performance, for example, it improves Recall@1 from 13.3% to 36.5% on E-VQA benchmark, 45.6% to 53.2% on InfoSeek, largely bridging the gap towards the ‘Google Lens’ upperbound.
@@
It is clear that our proposed EchoSight (w. reranking) has outperform the prior works by a significant margin, even approaching the upperbound results reported by original E-VQA Mensink et al. (2023) benchmark, where two giant models are adopted, i.e., ‘Google Lens’ for knowledge retrieval, and PaLM as answer generation.
@@ 3.5 Ablation Study
For all experiments in ablation study, we use the E-VQA dataset. On the retrieval side, we conduct the following ablation studies: (i) to compare different visual backbones in retrieval module, (ii) to study the impact of reranking scope and (iii) to investigate the importance of hard negative sampling. On final answer generation, we carry out ablation studies on: (i) the impact of different language models, (ii) to experiment the answer generator under oracle retrieval results.
@@
Impact of vision backbones. We assess the effect of different visual backbones on the retrieval stage, as detailed in Table 4. We compare the Vision Transformer (ViT) from EvaCLIP-8B Sun et al. (2024) with OpenAI’s CLIP-ViT-Large Radford et al. (2021). The EvaCLIP-8B’s ViT achieves a recall@20 of 48.8%, outperforming the CLIP-ViT-Large, which scored 32.2%. This substantial improvement is likely due to EvaCLIP-8B’s larger parameter size and more extensive training dataset, allowing it to develop more robust representations.
@@
Impact of reranking scope. The reranking scope refers to the number of candidates considered by the reranker module. Involving a higher reranking scope means calculating more embeddings during the reranking process. The reranking scope, which can be any number up to $k$, i.e., the total number of candidates returned by the retriever. As shown in Table 5, our reranker can consistently improve the results with increasing scope from Top-5 to Top-500, though it will significantly increase the computation cost, resulting in diminishing returns. Considering the balance of efficiency and quality, the scope of 20 candidate entries is used when reporting our final VQA accuracy on E-VQA and InfoSeek.
@@
Effect of oracle retrieval. Oracle retrieval indicates that the correct Wikipedia entry is always provided for generating the answer. As shown in Table 8, LLMs can almost flawlessly answer the question if oracle retrieval is provided.
@@ 4.1 Visual Question Answering
Knowledge-based VQA. The task involves questions that require information not present in the image. Pioneering datasets like OK-VQA Marino et al. (2019) and A-OKVQA Schwenk et al. (2022), which include questions needing knowledge beyond what is visually depicted, necessitate the integration of external world knowledge and commonsense reasoning. However, both datasets focus primarily on commonsense and general world knowledge, often neglecting more specialized or encyclopedic facts, and they do not provide external knowledge bases.
To fill this gap, datasets such as Encyclopedic VQA (E-VQA) Mensink et al. (2023) and InfoSeek Chen et al. (2023) have been developed. These datasets utilize Wikipedia as a knowledge base to provide detailed and specific information on various topics. E-VQA covers a wide range of topics like animals, plants, and landmarks, while InfoSeek focuses on info-seeking questions about various visual entities. These datasets require models to recognize visual entities and accurately retrieve and use relevant information from external sources Lerner et al. (2024); Caffagni et al. (2024); Lin et al. (2024).
@@ A.2 InfoSeek
Appendix B Qualitative Results
Figure 3: Qualitative results showing the effective performance of our multimodal reranking method. While the reference image and question is displayed at left side, the visual-only search results are showcased at row 1 of each example, and the multimodal reranking results are showcased at row 2. The target image is squared with orange box.
B.1 reranking results
Qualitative results of our EchoSight’s multimodal reranking are as shown in Figure 3.
B.2 VQA results

### v2

@@ Abstract
Knowledge-based Visual Question Answering (KVQA) tasks require answering questions about images using extensive background knowledge. Despite significant advancements, the large generative visual-language models often struggle with these tasks due to the limited integration of external knowledge. In this paper, we introduce EchoSight, a novel multimodal Retrieval-Augmented Generation (RAG) framework that enables to answer visual questions requiring fine-grained encyclopedic knowledge. To strive for high-performing retrieval, EchoSight first searches wiki articles by using visual-only information, subsequently, these candidate articles are further reranked according to their relevance to the combined text-image query. This approach significantly improves the integration of multimodal knowledge, leading to enhanced retrieval outcomes and more accurate VQA responses. Our experimental results on the Encyclopedic VQA and InfoSeek datasets demonstrate that EchoSight establishes new state-of-the-art results in knowledge-based VQA, achieving an accuracy of 41.8% on Encyclopedic VQA and 31.3% on InfoSeek.
@@
Yibin Yan           Weidi Xie School of Artificial Intelligence, Shanghai Jiao Tong University https://go2heart.github.io/echosight
@@ 1 Introduction
Despite the recent accomplishments in developing Visual-language Models (VLMs) Achiam et al. (2023); Gemini Team et al. (2023); Abdin et al. (2024); Liu et al. (2024), knowledge-based VQA remains challenging. This complexity primarily stems from two aspects. (i) Existing VLMs struggle to adequately encode all essential knowledge, due to its limited model capacity, and infrequent inclusion of encyclopedic, long-tail information in their training data Mensink et al. (2023). (ii) The visual component of the questions often provides limited help in addressing the queries, as establishing a meaningful connection between entity knowledge and visual attributes can be difficult. For example, an image of a church alone barely reveal information about its construction date.
@@
Overall, we present three contributions: First, we propose a multimodal retrieval-augmented generation framework, termed as EchoSight, that enables to answer visual questions that require fine-grained encyclopedic knowledge; Second, we adopt a retrieval-and-reranking scheme to improve retrieval performance, specifically, it first searches images with visual-only information, and then conduct a fine-grained multimodal reranking on the candidates; Third, we conduct thorough experiments on both Encyclopedic VQA Mensink et al. (2023) and InfoSeek Chen et al. (2023) benchmarks, EchoSight demonstrates state-of-the-art performance on both benchmarks, significantly outperforming existing VLMs or other retrieval-augmented architectures.
@@ 2.2 Retrieval and Reranking
On the candidates side, we break each of the wiki articles into sections, with each section prefixed by the article’s title, for example, $a_{i}=\{\text{sec}_{1}^{i},\text{sec}_{2}^{i},\dots,\text{sec}_{p}^{i}\}$, and further encode them with Q-Former’s text encoder. We initialize the Q-Former with BLIP-2’s weights and fine-tune all parameters except the visual encoder.
@@
$\text{sec}_{vl}=\operatorname*{arg\,max}_{\text{sec}\in a}\left(\alpha\cdot S_{v}^{\text{sec}}+(1-\alpha)\cdot S_{r}^{\text{sec}}\right),$
where $\text{sec}_{vl}$ refers to the highest-ranked entry section produced by the reranker, $\alpha$ is a weight parameter that balances the visual similarity score $S_{v}^{\text{sec}}$ and the reranking score $S_{r}^{\text{sec}}$. Note that, $S_{v}^{sec}$ is calculated in the visual-only search stage using the best-matched image from the wiki entry to which $sec$ belongs.
Reranker Training. Here, we implement hard negative sampling within a contrastive learning framework. Specifically, the negative samples are specifically selected from examples that are visually similar yet contextually distinct, i.e., the initial visual-only retrieval efforts were unsuccessful. With such training, the reranker is thus forced to select the most relevant articles for the multimodal queries, enhancing the overall accuracy and effectiveness of the system Robinson et al. (2021).
@@
$\mathcal{L}=-\log\frac{\exp(\max_{1\leq i\leq N_{q}}\text{sim}(z_{m}^{i},z_{s})/\mathcal{T})}{\sum_{j=1}^{N}\exp(\max_{1\leq i\leq N_{q}}\text{sim}(z_{m}^{i},z_{s}^{j})/\mathcal{T})},$
where $z_{s}$ denotes the positive section embedding, $N$ is the total number of samples including both positive and negatives and $\mathcal{T}$ is the temperature parameter that controls the smoothness of the softmax distribution.
@@ 3.3 Implementation Details
We adopt OneCycleLR Smith and Topin (2019) scheduler, with AdamW Loshchilov and Hutter (2018) optimizer. We use learning rate $10^{-4}$, batch size $6$, and the negative samples per example being $24$. For training the reranker module with 900K examples in Encyclopedic VQA, 150K steps require 40 hours on 1 Nvidia A100 (80G).
@@ 3.4 Results
On Retrieval. The experiment results for the retrieval tasks across different configurations are detailed in Table 1 and Table 2. The CLIP I-T setting involves using CLIP for cross-modal similarity search, from the reference image to the Wikipedia article. The articles are represented as CLIP embedding of their title and descriptions. The ‘Google Lens’ refers to the approach used in Encyclopedic VQA Mensink et al. (2023), where Google Lens indexes billions of images from the Internet, not limited to Wikipedia, to find and return the most closely matching images along with an entity prediction. The best corresponding knowledge base entry identified by Google Lens is considered as its retrieval results. Given its vast image index, which potentially includes the image from the test set and capability to associate images with relevant entities, Google’s retrieval can be viewed as a top performer in E-VQA retrieval.
From both tables, we can draw the observation that, our proposed reranking module has shown to significantly improve the retrieval performance, for example, it improves Recall@1 from 13.3% to 36.5% on E-VQA benchmark, 45.6% to 53.2% on InfoSeek, largely bridging the gap towards the ‘Google Lens’ top performer.
@@
It is clear that our proposed EchoSight (w. reranking) has outperform the prior works by a significant margin, even approaching the upperbound results reported by original E-VQA Mensink et al. (2023) benchmark, where two giant models are adopted, i.e., ‘Google Lens’ for knowledge retrieval, and PaLM for answer generation.
@@ 3.5 Ablation Study
For all ablation studies, we use the E-VQA dataset. On the retrieval side, we conduct the following ablations: (i) to compare different vision backbones in retrieval, (ii) to study the impact of reranking scope, and (iii) to investigate the importance of hard negative sampling. On final answer generation, we carry out ablation studies on: (i) the impact of various language models and (ii) to experiment with the answer generator under oracle settings.
@@
Impact of vision backbones. We assess the effect of different vision backbones on the retrieval stage, as detailed in Table 4. We compare the Vision Transformer (ViT) from EvaCLIP-8B Sun et al. (2024) with OpenAI’s CLIP-ViT-Large Radford et al. (2021). The EvaCLIP-8B’s ViT achieves a recall@20 of 48.8%, outperforming the CLIP-ViT-Large, which scored 32.2%. This substantial improvement is likely due to EvaCLIP-8B’s larger parameter size and more extensive training dataset, allowing it to develop more robust representations.
@@
Impact of reranking scope. The reranking scope refers to the number of candidates considered by the reranker module. Involving a higher reranking scope means calculating more embeddings during the reranking process. The reranking scope, which can be any number up to $k$, i.e., the total number of candidates returned by the retriever. As shown in Table 5, our reranker can consistently improve the results with increasing scope from Top-5 to Top-500. As the thoughput experiment showed in Table 6, considering the balance of efficiency and quality, the scope of 20 candidate entries is used when reporting our final VQA accuracy on E-VQA and InfoSeek.
@@
Scope Total Retrieval Time Throughput
Top 10 0.602 1.66
Top 20 1.171 0.85
Top 50 2.720 0.37
Top 100 5.082 0.20
Top 500 21.591 0.05
Table 6: Throughput Study. This study uses a NVIDIA A100 GPU with 80GB memory limiting the batch size to 1. For visual-only retrieval, we use the Faiss library with an exhaustive search. The throughput is calculated as the number of queries processed per second (QPS).
@@
Effect of oracle retrieval. Oracle retrieval indicates that the correct Wikipedia entry is always provided for generating the answer. As shown in Table 9, LLMs can almost answer the question if oracle retrieval is provided.
@@
     PaLM      KB Section      87.0
     Mistral-7B      KB Section      89.9
     LLaMA3-8B      KB Section      90.7
@@ 4.1 Visual Question Answering
Knowledge-based VQA. The task involves questions that require information not present in the image. Pioneering datasets like OK-VQA Marino et al. (2019) and A-OKVQA Schwenk et al. (2022), which include questions needing knowledge beyond what is visually depicted, necessitate the integration of external world knowledge and commonsense reasoning. Uni-modal knowledge bases like GS112K Luo et al. (2021) and Wiki21M Karpukhin et al. (2020) are adopted in prior works Lin et al. (2023); Lin and Byrne (2022); Gao et al. (2022); Luo et al. (2021, 2023). However, uni-modal knowledge bases are text-only, which limits their applicability in scenarios where visual context is paramount. To better utilize multimodal information, multiple previous attempts have been made Ding et al. (2022); Zhu et al. (2020); Wu et al. (2022); Chen et al. (2022).
Datasets such as Encyclopedic VQA (E-VQA) Mensink et al. (2023) and InfoSeek Chen et al. (2023) have been developed with multimodal knowledge bases. These datasets utilize Wikipedia as a multimodal knowledge base to provide detailed and specific information on various topics. E-VQA covers a wide range of topics like animals, plants, and landmarks, while InfoSeek focuses on info-seeking questions about various visual entities. These datasets require models to recognize visual entities and accurately retrieve and use relevant information from external sources Lerner et al. (2024); Caffagni et al. (2024); Lin et al. (2024).
@@ 5 Conclusion
Acknowledgements
This work is funded by National Key R&D Program of China (No.2022ZD0161400).
@@ A.2 InfoSeek
Appendix B Vision backbones
In addition to CLIP, there are other robust vision backbones available, such as Dino Caron et al. (2021); Oquab et al. (2023). Unlike CLIP, which employs a visual-language training method, Dino leverages a self-supervised, visual-focused training approach. To evaluate its performance, we benchmarked DinoV2 as our visual-only retriever, presenting the results in Tables 11 and 12. Despite observing a notable performance improvement in the Encyclopedic VQA task, there was a significant drop in performance on the InfoSeek task. Therefore, to maintain the consistency and overall performance of EchoSight, we have decided to continue using Eva-CLIP as our vision backbone.
Backbone Recall@K
K=1 K=5 K=10 K=20
Eva-CLIP
w/o. Reranking 13.3 31.3 41.0 48.8
w. Reranking 36.5 47.9 48.8 48.8
DINOv2
w/o. Reranking 17.3 38.6 46.0 51.4
w. Reranking 40.8 50.7 51.3 51.4
Table 11: DINOv2 comparison to CLIP in E-VQA.
Backbone Recall@K
K=1 K=5 K=10 K=20
Eva-CLIP
w/o. Reranking 45.6 67.1 73.0 77.9
w. Reranking 53.2 74.0 77.4 77.9
DINOv2
w/o. Reranking 34.7 53.4 60.0 64.5
w. Reranking 38.2 60.0 64.1 64.5
Table 12: DINOv2 comparison to CLIP in InfoSeek.